# FlightKooba: A Fast Interpretable FTP Model

### v1

@@ 5.3 Accuracy
Due to the large number and high dimensionality of the datasets, it is not possible to provide predictions for all features in this section. Therefore, only one or two features from each dataset are shown in the test set. (Figures 6 and 7)
Figure 6: Selected prediction results for the three models on the test set for each dataset (a).
Figure 7: Selected prediction results for the three models on the test set for each dataset (b).
@@
5.4 Summary And Analysis
In this section, we tested the performance of the FlightKooba, Mamba, and xLSTM modules on 11 datasets. Specifically, the performance of the three models was compared in terms of convergence speed, accuracy, training time, and memory usage. The experimental results show that the FlightKooba module exhibits significant advantages in terms of training time, memory usage, and total number of parameters. Furthermore, the model accuracy achieved on the physical system dataset is on par with Mamba and xLSTM. This proves the feasibility of this calculation method, which makes it possible to introduce state space equations based on automatic control principles to construct and calculate Koopman operators.
Thanks to the SSM structure, Mamba and FlightKooba required less time for training. Compared to Mamba, FlightKooba uses state space equations to link the HIPPO method with Koopman theory, which is equivalent to determining the values of matrices $A$, $B$, and $C$ in the original SSM structure, further reducing the number of parameters that need to be calculated in the model.
However, FlightKooba performs poorly in certain situations. The following reasons are speculated: (a) The control variable $u$ selected is not the precise control variable for the system, leading to a certain degree of deviation in the prediction results; (b) The approximation errors in the derivation process affect the model to some extent; (c) The polynomial selected by the HIPPO method is a Legendre polynomial, which involves factorial calculations in the module,directly affecting the range of polynomial terms (the float type can only store factorial values up to approximately 32).
@@ 6 Conclusion And Future Works
In this paper, we propose a modeling and control framework based on the HIPPO method, Koopman theory, and state space equations from cybernetics: FlightKooba. This greatly improves the interpretability of the model and significantly reduces the number of trainable parameters. To evaluate the performance of FlightKooba, we tested and compared three models, FlightKooba, Mamba, and xLSTM, on multiple different datasets, including CAFUC and Lorenz. Experiments have demonstrated the superiority of the FlightKooba modeling method in terms of time and memory consumption (training time comparable to the Mamba module without CUDA-level acceleration; memory reduced by more than 50% on most datasets, with a tenfold reduction in the number of parameters), effectively completing the FTP task. This provides a new method for the rapid calculation of Koopman operators, opening up new possibilities for the combination of time series prediction and control.
Similarly, in our experiment, we identified shortcomings that FlightKooba will need to address in the future:
@@
    Other more suitable polynomials can be found to replace Legendre polynomials to fit the observed function, thereby eliminating the influence of factorials on the model.
@@
    A unique physical acceleration algorithm can be constructed to further reduce the time required for model training and prediction.
@@
    Modify the structure of FlightKooba to achieve higher accuracy in different physical systems (e.g., control variable $u$ only affects the state vector, and interactions between different features are not reflected).
And applying the modeling framework based on FlightKooba to high-reliability intelligent aviation systems to support application scenarios for intelligent aviation management.
Appendix A Details of The HIPPO Algorithm
When we use models similar to RNN networks (including SSM) for learning, the output of each iteration is only related to the input and hidden layer state of the current step. The disadvantage is obvious: as time progresses, the historical information recorded by the model gradually decreases.
The HIPPO method was proposed to store data for a period of time using a finite-dimensional vector: the data for this period of time is regarded as a function $\gamma(t)$, which is approximated by a polynomial:
$\gamma(t)=\sum_{i=0}^{n}c_{n}g_{n}(t)$ (36)
In this formula, $g_{n}$ is a basis function with respect to time $t$, and $c_{n}$ is the coefficient preceding each basis function. In this formula, we do not know the values of the coefficients $c_{n}$, which need to be calculated.For convenience of calculation, we specify that the basis functions $g_{n}$ must satisfy the properties of standard orthogonal bases (the HIPPO method uses many types of orthogonal bases, and here we choose the normalized Legendre polynomials).
Referring to the idea of finding the coefficient $c_{n}$ using Fourier series, we multiply both sides of 36 by the conjugate $g_{n}$ to obtain the formula for calculating the coefficient $c_{n}$ (the domain of the normalized Legendre polynomial $g_{n}$ is $[-1,1]$):
[…]
In order to make the above equation fit a more general time $t\in[0,T]$, find a mapping $s\mapsto t\leq T(s)$ from $[a,b]$ to $[0,T]$ and then derive the coefficient calculation formula (In the HIPPO method, the mapping $t\leq T(s)=(s+1)\omega/2+T-\omega$ used by $LegT$ and the mapping used by $LegS$ are $t\leq T(s)=(s+1)T/2$):
[…]
Then, differentiate both sides of 38 and use integration by parts to obtain:
$\frac{d}{dT}c_{n}(T)$ $=\gamma(t\leq T(s))\left(\frac{\partial t\leq T(s)}{\partial T}/\frac{\partial t\leq T(s)}{\partial s}\right)\Big{|}_{s=b}^{s=a}g_{n}(s)$
$\qquad-\int_{a}^{b}\gamma(t\leq T(s))d\left[\left(\frac{\partial t\leq T(s)}{\partial T}/\frac{\partial t\leq T(s)}{\partial s}\right)g_{n}(s)\,\right]$ (39)
The normalized Legendre polynomials $g_{n}$ and the Legendre polynomials $p_{n}$ are related by the following equation:
$g_{n}(t)=\sqrt{\frac{2n+1}{2}}p_{n}(t)$ (40)
The Legendre polynomials and their derivatives are related by a recursive formula:
$\frac{d}{dt}p_{n}(t)=np_{n-1}(t)$ (41)
The idea behind $LegT$ is basically the same as that behind $LegS$. Let us derive LegT below: Introduce a mapping $t\leq T(s)=(s+1)\omega/2+T-\omega$, which maps $[-1,1]$ to $[T-\omega,T]$ (meaning that we only care about the history of length $\omega$), and then substitute it into equation 39 to obtain:
$\frac{d}{dT}c_{n}(T)$ $=\gamma(t\leq T(s))\left(\frac{\partial t\leq T(s)}{\partial T}/\frac{\partial t\leq T(s)}{\partial s}\right)g_{n}(s)\Big{|}_{s=-1}^{s=1}$
$\qquad-\int_{-1}^{1}\gamma(t\leq T(s))d\left[\left(\frac{\partial t\leq T(s)}{\partial T}/\frac{\partial t\leq T(s)}{\partial s}\right)g_{n}(s)\right]$
$=\sqrt{2(2n+1)}\left[\frac{1}{\omega}\gamma(T)-(-1)^{n}\frac{1}{\omega}\gamma(T-\omega)\right]$
$\qquad-\int_{-1}^{1}\gamma(t\leq T(s))d\left[\left(\frac{\partial t\leq T(s)}{\partial T}/\frac{\partial t\leq T(s)}{\partial s}\right)g_{n}(s)\right]$
$=\sqrt{2(2n+1)}\left[\frac{1}{\omega}\gamma(T)-(-1)^{n}\frac{1}{\omega}\gamma(T-\omega)\right]$
[…]
$=\sqrt{2(2n+1)}\left[\frac{1}{\omega}\gamma(T)-(-1)^{n}\frac{1}{\omega}\gamma(T-\omega)\right]$
[…]
To further solve this problem, another recursive formula from Legendre polynomials is used here:
$p_{n+1}^{\prime}(t)-p_{n-1}^{\prime}(t)=(2n+1)p_{n}(t)$ (43)
Expanding the above equation, we obtain:
$p_{n+1}^{\prime}(t)$ $=(2n+1)p_{n}(t)+(2n-3)p_{n-2}(t)+...$
$=\sum_{k=0}^{n}(2k+1)\tau_{n-k}p_{k}(t)$ (44)
Among them, when $k$ is even, $\tau_{n-k}=1$; otherwise, $\tau_{n-k}=0$.
Therefore, equation 42 can be further derived as follows:
$\frac{d}{dT}c_{n}(T)$ $=\sqrt{2(2n+1)}\left[\frac{1}{\omega}\gamma(T)-(-1)^{n}\frac{1}{\omega}\gamma(T-\omega)\right]$
[…]
$=\sqrt{2(2n+1)}\left[\frac{1}{\omega}\gamma(T)-(-1)^{n}\frac{1}{\omega}\gamma(T-\omega)\right]$
$\qquad-\frac{2}{\omega}\int_{-1}^{1}\gamma(t\leq T(s))\sqrt{\frac{2n+1}{2}}\sum_{k=0}^{n}(2k+1)\tau_{n-k-1}p_{k}(s)ds$
$=\frac{\sqrt{2(2n+1)}}{\omega}[\gamma(T)-(-1)^{n}\gamma(T-\omega)]$
$\qquad-\frac{2}{\omega}\int_{-1}^{1}\gamma(t\leq T(s))\sqrt{\frac{2n+1}{2}}\sum_{k=0}^{n-1}\sqrt{2(2k+1)}\tau_{n-k-1}g_{k}(s)ds$
$=\frac{\sqrt{2(2n+1)}}{\omega}[\gamma(T)-(-1)^{n}\gamma(T-\omega)]$
$\qquad-\frac{\sqrt{2(2n+1)}}{\omega}\int_{-1}^{1}\gamma(t\leq T(s))\sum_{k=0}^{n-1}\sqrt{2(2k+1)}\tau_{n-k-1}g_{k}(s)ds$
$=\frac{\sqrt{2(2n+1)}}{\omega}[\gamma(T)-(-1)^{n}\gamma(T-\omega)]$
$\qquad-\frac{\sqrt{2(2n+1)}}{\omega}\sum_{k=0}^{n-1}\sqrt{2(2k+1)}\tau_{n-k-1}c_{k}(T)$ (45)
From 36 we obtain:
$\gamma(t\leq T(s))\approx\sum_{n=0}^{N}c_{n}g_{n}\left(s\right)$ (46)
Let $s=-1$, we obtain:
$\gamma(T-\omega)\approx\sum_{k=0}^{N}c_{k}g_{k}\left(-1\right)=\sum_{k=0}^{N}(-1)^{k}c_{k}(T)\sqrt{\frac{2k+1}{2}}$ (47)
Substituting 47 into 45, we obtain:
$\frac{d}{dT}c_{n}(T)$ $=\frac{\sqrt{2(2n+1)}}{\omega}[\gamma(T)-(-1)^{n}\gamma(T-\omega)]$
$\qquad-\frac{\sqrt{2(2n+1)}}{\omega}\sum_{k=0}^{n-1}\sqrt{2(2k+1)}\tau_{n-k-1}c_{k}(T)$
$=\frac{\sqrt{2(2n+1)}}{\omega}\gamma(T)-\frac{\sqrt{(2n+1)}}{\omega}\sum_{k=0}^{N}(-1)^{n-k}c_{k}(T)\sqrt{2k+1}$
$\qquad-\frac{\sqrt{2n+1}}{\omega}\sum_{k=0}^{n-1}\sqrt{(2k+1)}2\tau_{n-k-1}c_{k}(T)$
$=\frac{\sqrt{2(2n+1)}}{\omega}\gamma(T)-\frac{\sqrt{(2n+1)}}{\omega}\sum_{k=n}^{N}(-1)^{n-k}c_{k}(T)\sqrt{2k+1}$
$\qquad-\frac{\sqrt{2n+1}}{\omega}\sum_{k=0}^{n-1}\sqrt{(2k+1)}(2\tau_{n-k-1}+(-1)^{n-k})c_{k}(T)$ (48)
Among these, $2\tau_{n-k-1}+(-1)^{n-k}$ represents the number of terms used in the polynomial approximation of $\gamma(t)$. Therefore, we can choose n to be an even number, i.e., use an even number of terms for the approximation. Based on this, n-k-1 and n-k must be one odd and one even number. When n-k is even, k is even, then $02\tau_{n-k-1}=$ 0 and $(-1)^{n-k}=1$; when n-k is odd, k is odd, then $22\tau_{n-k-1}=$ 2 and $(-1)^{n-k}=-1$. That is, $2\tau_{n-k-1}+(-1)^{n-k}\equiv 1$. Therefore, Equation 48 can be simplified to:
$\frac{d}{dT}c_{n}(T)$ $=\frac{\sqrt{2(2n+1)}}{\omega}\gamma(T)-\frac{\sqrt{(2n+1)}}{\omega}\sum_{k=n}^{N}(-1)^{n-k}c_{k}(T)\sqrt{2k+1}$
$\qquad-\frac{\sqrt{2n+1}}{\omega}\sum_{k=0}^{n-1}\sqrt{(2k+1)}c_{k}(T)$ (49)
The above equation can be converted into matrix form:
@@ Appendix A Details of The HIPPO Algorithm
Among them, $c(t)=[c_{0}(T),c_{1}(T),\ldots,c_{N}(T)]$, and the expressions of matrices $N$ and $M$ are as follows:
$N_{n,k}=-\frac{1}{\omega}\begin{cases}\sqrt{(2n+1)(2k+1)},&k<n\\ (-1)^{n-k}\sqrt{(2n+1)(2k+1)},&k\geq n\end{cases}$ (51)
$M_{n}=\frac{1}{\omega}\sqrt{2(2n+1)}$ (52)
Introducing the scaling factor $\lambda_{n}=\frac{2}{\sqrt{2n+1}}$, such that $c_{n}(T)=\lambda_{n}\tilde{c}_{n}(T)$, the expressions for matrices $N$ and $M$ become:
$N_{n,k}=-\frac{1}{\omega}\begin{cases}2n+1,&k<n\\ (-1)^{n-k}(2n+1),&k\geq n\end{cases}$ (53)
$M_{n}=\frac{1}{\omega}(2n+1)$ (54)
Following the same line of reasoning, introducing the mapping $t\leq T(s)=(s+1)T/2$ and the scaling factor $\lambda_{n}=\sqrt{2}$, we obtain the expressions for the LegS method matrices N and M:
$N_{n,k}=\begin{cases}\sqrt{(2n+1)(2k+1)},&k<n\\ n+1,&k=n\\ 0,&k>n\end{cases}$ (55)
$M_{n}=\sqrt{2(2n+1)}$ (56)
Appendix B Multi-step Iteration
@@
Then, we move on to the discretization on the right side of the equation. Since the data we obtain is discrete, we assume that $\gamma(t)$ is a fixed value between $[t_{k},t_{k+1}]$, so we can discretize $\gamma(t)$ as $\gamma_{k}=\gamma(t_{k})$. There are two ways to discretize $c(t)$. One assumes that $c_{t_{k}}=c(t_{k})$ in the interval $[t_{k},t_{k+1}]$, and the other assumes that $c_{t_{k+1}}=c(t_{k})$, i.e., the two values at the endpoints of the interval. If we consider $c_{t_{k}}=c(t_{k})$, we obtain:
@@
With the above two equations 59 and 61, a third method for estimating the value of $c(t_{k})$ emerges, which assumes that:
@@
$=R\bar{M}\gamma(t_{k})$ (72)
where $R=(\overline{N}^{k-1},\overline{N}^{k-2},\ldots,\overline{N},I)$,$\overline{M}=\Delta\mathrm{t}\left(I-\frac{\Delta t}{2}N\right)^{-1}M$. $\gamma(t_{k})=(\gamma_{1},\gamma_{2},\ldots,\gamma_{k})^{T}$ denotes the observed values of $\gamma(t)$ from $\gamma_{k}$ backward to the discrete values $k-1$ to the discrete value $k$th (a total of values $k$, i.e. a total of $k$ iterations.
@@
It can also be written in matrix form similar to 72:
@@
Therefore, during the actual iteration, it is only necessary to calculate the $k-1$th power of the matrix once and store the result, which can then be used until the iteration ends, thereby reducing the time complexity of the algorithm.

### v2

@@ 5.2 Accuracy
The results of the accuracy test for each model in the test set are shown in Figures 4 and 5. Due to the large dataset size and feature dimensions, only a subset of results is presented here for demonstration purposes.
Figure 4: Prediction results for some dimensions on the test set.
Figure 5: Prediction results for some dimensions on the test set.
@@
5.3 Analysis of Model Limitations
Although comprehensive experiments across 11 datasets (as shown in Tables 1 and 2) highlight FlightKooba’s indisputable computational efficiency—achieving the fastest training speed, lowest memory consumption, and parameter counts several orders of magnitude smaller than advanced baselines like Mamba, xLSTM, and Koopa—our analysis reveals a unique bimodal performance pattern regarding prediction accuracy. Specifically, on test sets, FlightKooba either performs exceptionally well, nearly matching the accuracy of complex models, or fails to capture meaningful dynamics, with few intermediate cases. This section delves into this phenomenon and traces it back to the theoretical foundations of our model.
We observe three key behavioral characteristics. First, on datasets dominated by high-frequency components like the sixth feature in CAFUC and ETTh2, FlightKooba’s predictive performance sharply deteriorates. The model’s output either collapses into a nearly constant line approximating the signal mean or learns only the locations of peaks and troughs with significantly reduced amplitude, indicating its inability to capture underlying dynamics (as shown in 6). Second, for periodic signals successfully tracked by the model (e.g., on traffic and electricity datasets), we note persistent phase lags where predicted peaks and troughs trail slightly behind actual values. Third, predicted amplitudes are frequently attenuated, causing forecast curves to appear “flatter” than the true signal.
Figure 6: Illustration of model limitations on high-frequency datasets. Left: Noise signal from CAFUC (Feature 6), where predictions collapse to the signal mean. Right: Model fails to capture dynamics in ETTh2 (Feature 1), consistent with its low-pass filtering characteristics.
Initially, these observations cast doubt on the validity of our theoretical derivations. However, after thorough examination and validation, we confirmed that these behaviors are not flaws but direct and predictable consequences of our model’s core design. FlightKooba is fundamentally based on the HiPPO framework, which approximates time series using a set of Legendre polynomial bases. From a signal processing perspective, this entire mechanism structurally equates to a powerful low-pass filter. The inherent purpose of a low-pass filter is to suppress high-frequency noise while capturing the primary, smooth long-term trends and periodic components within a signal. This intrinsic property defines both the model’s strengths and limitations: its strength lies in effectively filtering out noise to reveal the core smooth skeleton of signal dynamics; its limitation stems from its structural design to inherently ignore high-frequency fluctuations and abrupt transients. This theoretical perspective provides a clear explanation for our experimental results. For signals predominantly composed of high-frequency, noise-like components (such as the sixth feature in CAFUC), the model effectively filters out all dynamic information, leaving only the statistical mean. This also accounts for the observed phase lag and amplitude attenuation—both typical phenomena when applying a low-pass filter to dynamic signals.
Based on this, we can define ideal and unsuitable application scenarios for FlightKooba. Ideal Application Scenarios:
  • 1.
    Systems with clear physical underlying structures and fundamental principles (e.g., weather system evolution).
  • 2.
    Signals with strong periodic or seasonal components (e.g., traffic flow, power load, epidemic spread).
  • 3.
    Applications where computational resources and real-time performance are critical constraints, focusing on core trends rather than fine-grained fluctuations.
Unsuitable Application Scenarios:
  • 1.
    Signals with high-frequency components where random noise itself contains significant information.
  • 2.
    Chaotic system predictions (e.g., Lorenz systems) where minor perturbations cause substantial long-term divergence.
  • 3.
    Financial time series (e.g., stock prices) exhibiting behavior closer to random walks.
In summary, our findings demonstrate that FlightKooba successfully achieves a highly favorable trade-off between computational cost and prediction performance within its intended application domains, validating its viability as a lightweight and resource-efficient solution. Future work can extend this research in two primary directions. First, exploring FlightKooba as an efficient “backbone extractor” may involve adding a lightweight “high-frequency compensation module” to cover a broader range of signal types. Second, investigating the replacement of the Legendre polynomial basis with alternative functional bases (e.g., wavelets) better suited to specific signal characteristics could enhance the model’s flexibility.
@@ 6 Conclusion And Future Works
In this paper, we introduce FlightKooba, a novel time series modeling approach prioritizing computational efficiency and interpretability. Our comprehensive evaluation across 11 datasets demonstrates its significant breakthrough in this domain. Compared to established modules like Mamba and xLSTM, as well as state-of-the-art models such as Koopa and flightbert++, FlightKooba consistently achieves superior performance in training time, memory usage, and parameter efficiency. This exceptional efficiency stems directly from its fundamental design: FlightKooba avoids extensive parameter training by analytically deriving its state space matrices (A, B, and C) through the synergistic integration of HiPPO and Koopman theory. This approach not only drastically reduces model complexity but also validates the feasibility of systematically constructing Koopman operators from control theory principles, paving a new path for lightweight, interpretable models.
However, as discussed in our analysis, the model’s strengths are intrinsically linked to its limitations. We identify three key areas for future research aimed at broadening its applicability:
@@
    Explicit control variable integration: Due to dataset constraints, our benchmarking utilized proxy control variables (i.e., the last $n$ dimensional features per dataset). A crucial next step is applying FlightKooba to systems with explicit control inputs ($u$), which promises to unlock its full predictive potential and validate its design within a genuine control theory context.
@@
    Quantification of approximation errors: The model’s mathematical derivation relies on approximations inherent to the HiPPO approach and Koopman theory. Systematically quantifying the impact of these approximation errors on final prediction accuracy represents a vital research direction for enhancing model reliability.
@@
    Enhancing Numerical Stability and Expressiveness: While effective, the use of Legendre polynomials introduces factorial computations that limit model complexity due to numerical precision constraints. Future work should explore alternative polynomial bases or structural optimization modifications to enhance numerical stability and extend the model’s ability to capture more complex dynamics without sacrificing efficiency.
In summary, we position FlightKooba not as a universal replacement for all time series models, but as a powerful, computationally superior alternative for addressing a specific, critical class of problems. While trade-offs in predictive accuracy exist—particularly for high-frequency signals—its performance in scenarios where efficiency and interpretability are paramount validates its value as a significant contribution to the field.
Appendix A Multi-step Iteration
In the HIPPO algorithm, we obtain the following formula:
@@ Appendix A Multi-step Iteration
Then, we move on to the discretization on the right side of the equation. Since the data we obtain is discrete, we assume that $\gamma(t)$ is a fixed value between $t_{k},t_{k+1}]$, so we can discretize $\gamma(t)$ as $\gamma_{k}=\gamma(t_{k})$. There are two ways to discretize $c(t)$. One assumes that $c_{t_{k}}=c(t_{k})$ in the interval $t_{k},t_{k+1}]$, and the other assumes that $c_{t_{k+1}}=c(t_{k})$, i.e., the two values at the endpoints of the interval. If we consider $c_{t_{k}}=c(t_{k})$, we obtain:
@@
With the above two equations 39 and 41, a third method for estimating the value of $c(t_{k})$ emerges, which assumes that:
@@
$=R\overline{M}\gamma(t_{k})$ (52)
where $R=(\overline{N}^{k-1},\overline{N}^{k-2},\ldots,\overline{N},I)$$\overline{M}=\Delta\mathrm{t}\left(I-\frac{\Delta t}{2}N\right)^{-1}M$. $\gamma(t_{k})=(\gamma_{1},\gamma_{2},\ldots,\gamma_{k})^{T}$ denotes the observed values of $\gamma(t)$ from $\gamma_{k}$ backward to the discrete values k1k-1 to the discrete value $k$th (a total of values $k$, i.e. a total of $k$ iterations.
@@
It can also be written in matrix form similar to 52:
@@
Therefore, during the actual iteration, it is only necessary to calculate the k1k-1th power of the matrix once and store the result, which can then be used until the iteration ends, thereby reducing the time complexity of the algorithm.